OpenStudy (anonymous):

y=x^3 - 4 find the inverse function

OpenStudy (anonymous):

solve for x x = (y-4)^(1/3) so f^-1(x) = (x-4)^(1/3)

OpenStudy (anonymous):

i meant f^-1(x) = (x+4)^(1/3) *plus sign*